OpenStudy (anonymous):

What is a theoretical probability of rolling a sum of 10 on one roll of two standard number cubes?

OpenStudy (anonymous):

First count how many ways there are to get each value from 2-12. Outcome List of Combinations Total 2 1+1 1 3 1+2, 2+1 2 4 1+3, 2+2, 3+1 3 5 1+4, 2+3, 3+2, 4+1 4 6 1+5, 2+4, 3+3, 4+2, 5+1 5 7 1+6, 2+5, 3+4, 4+3, 5+2, 6+1 6 8 2+6, 3+5, 4+4, 5+3, 6+2 5 9 3+6, 4+5, 5+4, 6+3 4 10 4+6, 5+5, 6+4 3 <----- 11 5+6, 6+5 2 12 6+6 1 There are 36 possibilities (6x6) Outcome Probability 2 1/36 = 2.78% 3 2/36 = 5.56% 4 3/36 = 8.33% 5 4/36 = 11.11% 6 5/36 = 13.89% 7 6/36 = 16.67% 8 5/36 = 13.89% 9 4/36 = 11.11% 10 3/36 = 8.33% <----- 11 2/36 = 5.56% 12 1/36 = 2.78% 8 and 1/3 is your answer